# Large language models can be zero-shot anomaly detectors for time series?

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated an outstanding ability to learn natural language tasks implicitly, whether the task is reading comprehension, text summary, translation, etc. Radford et al. (2019); Brown et al. (2020); Sanh et al. (2022); Chowdhery et al. (2023); Wei et al. (2022). Moreover, LLMs have shown tremendous promise in formal language generation, such as code generation and synthesis Xu et al. (2022a); Austin et al. (2021); Chen et al. (2021) and in production beyond textual output, such as generating images and videos from natural language descriptions Saharia et al. (2022); Koh et al. (2023). Constantly testing these models on new tasks and data modalities allows us to push the boundaries of LLMs and discover their value. In this paper, we present a thorough study of LLMs used for the challenging task of anomaly detection from time series, asking the question Can LLMs become anomaly detectors for time series data? Here, LLMs are exposed to a new data type – time series – and are tasked with detection – different from a classic classification task, which they are known to excel at Howard and Ruder (2018).
@@
Figure 1: (left) Model F1 Score performance, compared to a moving average baseline. Each category represents a collection of models that fall under that group. For classic models, we consider ARIMA and Matrix Profiling; for Deep Learning (DL), we utilize AER and LSTM DT; for transformer anomaly detection models, we look at Anomaly Transformer; lastly, for the commercial category, we compare to MS Azure. (right) Illustration of Mistral forecasts on E-2 signal from the SMAP dataset. The deviation between the original signal and the forecasted one can help identify anomalous regions.
@@
Recently, Gruver et al. (2023) posited that large language models have an inherent auto-regressive feature which allows them to be effective forecasters. In the study, the authors fed a pretrained LLM a string representation of a time series sequence. The LLM then generated the next expected values, treating time series forecasting as a next-token prediction task. A follow-up question arises: Does LLMs’ auto-regressive nature allow them to take on more complex tasks, such as anomaly detection? State-of-the-art time series anomaly detection models using deep learning typically include a forecasting model as one of the steps in their process Hundman et al. (2018).
@@
Time series anomaly detection is commonly required in day-to-day industry operations. Identifying unexpected patterns that deviate from the norm can be a cumbersome task, especially when a massive amount of signals need to be analyzed. If LLMs are genuine anomaly detectors and can be employed directly in zero-shot (without any additional training), then they can serve as an off-the-shelf anomaly detector for users, lifting a considerable amount of this burden. Considering that training deep learning models is time-consuming, skipping this phase can increase the efficiency of anomaly detection.
@@
In this paper, we present SigLLM, a framework for finding anomalies in time series data using LLMs with current interaction support for models hosted by OpenAI  [^1] https://platform.openai.com/docs/models and HuggingFace  [^2] https://huggingface.co/models. Our framework includes a signal-to-text representation component to convert time series data into LLM-ready input. Moreover, we present two distinct approaches in investigating our main question. First, Prompter is a simple and direct prompting method to elicit LLMs to identify parts of a sequence it thinks are anomalous. Second, Detector leverages LLMs’ ability to forecast time series in order to find anomalies using the residual between the original signal and the forecasted one.
@@
Our findings, captured in Figure 1(left), show that LLMs improve on a simple moving average baseline. Moreover, they outperform transformer-based models such as Anomaly Transformer Xu et al. (2022b). However, there is still a gap between classic and deep learning approaches and LLMs. Furthermore, between our two approaches, Detector is superior to Prompter with an improvement of 135% in F1 Score, as the latter suffers from false positives. We highlight the potential of the Detector approach in Figure 1(right), which showcases an example of an LLM forecast. We can clearly see that the LLM forecast is substantially different from the original signal; this difference is attributed to the presence of anomalies.
@@
  • Propose a new application for LLMs–anomaly detection–and study their efficacy and efficiency for this task. We formalize a new task to present to large language models; namely, time series anomaly detection in zero-shot.
  • Present the SigLLM framework with a time-series-to-text representation module and two novel methodologies for solving this task. We present SigLLM with a module to convert time series data into language-model-ready input through a series of reversible transformations. Moreover, we propose two distinct approaches for solving the problem: the Prompter and Detector pipelines. As of writing this paper, our framework integrates propriety models such as GPT by OpenAI and open models provided on the HuggingFace transformers package.
  • Provide a comprehensive and thorough evaluation of how well LLMs perform this task. We conduct our experiments on two prominent LLM models: GPT-3.5-turbo and Mistral-7B-Instruct-v0.2, and 11 datasets. We show that LLMs are able to find anomalies with an average F1 score of 0.525. Moreover, we compare SigLLM methods to 10 other existing methods including state-of-the-art models such as AER Wong et al. (2022).
  • Publish an open source software. Our code and datasets are publicly available on github: https://github.com/ (removed to preserve anonymity).
@@ 2 Background and Related Work
Anomaly Detection Pipelines. Unsupervised machine learning-based anomaly detection pipelines generally follow the same sequence of steps, which roughly consist of pre-processing, modeling, and post-processing abstractions as presented in Figure 2 Alnegheimish et al. (2022). Pre-processing operations include scaling the time series into a specific range, while post-processing includes computing the discrepancies between two sequences. A wide variety of models can be trained to learn the features of input data, including Long-Short-Term-Memory models (LSTMs) Hundman et al. (2018), AutoEncoders (AE) Malhotra et al. (2016), Variational AutoEncoders (VAE) Park et al. (2018), Generative Adversarial Networks (GANs) Geiger et al. (2020), and Transformers Xu et al. (2022b); Tuli et al. (2022). These models are performative on existing benchmarks such as Alnegheimish et al. (2024), surpassing the performance of statistical approaches such as ARIMA Box and Pierce (1970); Pena et al. (2013).
@@
Transformers for Time Series. Transformers can be directly used to reconstruct time series Tuli et al. (2022). Moreover, the attention mechanism can be leveraged to find anomalous sequences Xu et al. (2022b). Recently, more work has emerged adopting transformer-based models for forecasting purposes by pretraining large transformer models on a large corpus of time series data. ForecastPFN Dooley et al. (2023) pre-trains a basic encoder-decoder transformer with one multi-head attention layer and two feedforward layers on a synthetically generated time series dataset. Similarly, TimeGPT was pretrained on a large collection of publicly available time series datasets. Lag-Llama Rasul et al. (2023) is a decoder-only Llama-2 model pretrained on a large corpus of real time series data from diverse domains. Moreover, Chronos Ansari et al. (2024) adopts a T5 architecture and parses time series data into text to pretrain their model. Most of these models were developed with the objective of creating a time series foundation model for time series forecasting.
@@
LLMs for Time Series. Over the past several months, we’ve seen considerable efforts toward LLM utilization for time series data. Given the parallels between predicting the next word in a sentence and predicting the next value in a time series, most of these efforts have focused on time series forecasting. One notable effort is LLMTime where Gruver et al. (2023) employ GPT Brown et al. (2020), and Llama-2 Touvron et al. (2023) models to forecast time series data. PromptCast Xue and Salim (2023) is a related work that translates a forecasting problem into a prompt, transforming forecasting into a question-answering task.
@@
Figure 3: Visualizing the output of large language models (GPT and Mistral) under different variations of the transformation process. Each row depicts the exchange-2_cpm_results signal from the AdEx dataset, where the x-axis shows the timestamp and the y-axis is the signal value. The first row indicates the ground truth anomalies present in the time series (highlighted in green). The remaining rows indicate whether scaling and inserting space between digits has occurred during the conversion from signal to text. The gray intervals highlight the anomalies detected under these conditions, thus, we would like to maximize the overlap between the green and gray intervals. Overall we find that “scaling + space” is the configuration that yields a better output for GPT; and “scaling + no space” is better for Mistral.
@@
Our Work. In this paper, we work strictly with LLMs that were pre-trained on text, particularly a proprietary model using GPT-3.5 Brown et al. (2020) and an open source model using Mistral Jiang et al. (2023). Our main objective is to determine whether LLMs have the ability to directly uncover anomalies in time series data. Referring back to Figure 2, our methodology focuses on step 2 onwards, primarily the inference phase. To our knowledge, there is no other work that utilizes large language models as zero-shot anomaly detectors for time series data. We explore two avenues for accomplishing this task: (a) Through the paradigm of prompt engineering; (b) By leveraging LLMs’ ability to forecast time series in zero-shot without any additional data or fine-tuning.
@@ 3 Time Series Representation
Time series data can be stored in many different forms. In this paper, we define a univariate time series as $\mathbf{X}=(x_{1},x_{2},\dots,x_{T})$, where $x_{t}\in\mathbb{Z}_{\geq 0}$ is the value at time step $t$, and $T$ is the length of the series. To make a time series LLM-ready, we transform the univariate time series $\mathbf{X}$ into a sequence of values that is tokenized. We follow a sequence of reversible steps, including scaling, quantization, processing the time series into segments using rolling windows, and lastly tokenizing each window. We detail these steps below.
@@
Figure 4: Anomaly detection methods in the SigLLM framework. (a) Prompter: a prompt engineering approach to elicit large language models in identifying parts of the input which are anomalies. (b) Detector: a forecasting approach to use large language models as forecasting methods, Detector then finds the discrepencies between the original and forecasted signal as indication for the presence of anomalies.
Other scaling methods, such as min-max scaling, can be utilized to achieve the same goal. However, reducing the set of possible values to a smaller range (e.g. [0, 1]), may cause loss of information in the quantization step. On the other hand, increasing the range will result in more digits to tokenize. With our approach, we simply shift the range of the signal values, which allows us to reduce the number of individual digits that need to be tokenized and maintains the original gaps between pairs of entries. Moreover, by projecting the values into a non-negative range, we eliminate the need for sign indicator “-/+” and save an additional token.
Quantization. Unlike the finite set of vocabulary words used to train LLMs (32k vocab tokens for Mistral [^3] The exact vocabulary size for GPT-3.5-turbo is not shared by OpenAI., the set of scaled time series values $x_{s_{t}}$ is infinite, and cannot be processed by language models. Therefore, quantization is generally done to prepare time series for LLMs Ansari et al. (2024); Gruver et al. (2023). We use the rounding method, as proposed in in Gruver et al. (2023). Specifically, because in some cases the number of decimal digits are redundant given a fixed precision, we round up each value to a predetermined number of decimals and subsequently scale to an integer format to avoid wasting tokens on the decimal point. Hence, the input time series becomes $\mathbf{X}_{q}=(x_{q_{1}},x_{q_{2}},\dots,x_{q_{T}})$, where $x_{q_{t}}\in\mathbb{Z}_{\geq 0}$. Below is an example of this operation :
@@
Overall, we use 2 mapping functions: the scaling function noted $\mathcal{E}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ and the quantization function noted $\mathcal{Q}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{Z}_{\geq 0}$. Since both mapping functions are reversible up to a certain number of precision digits, we can always reconstruct the input time series: $\mathcal{E}^{-1}\left(\mathcal{Q}^{-1}\left(x_{q_{t}}\right)\right)\approx x_{t}$
Rolling windows. Because there is an upper limit on the context length input to LLMs (e.g., Mistral has an upper limit of 32k tokens and GPT-3.5-turbo has a limit of 16k tokens), and constraints on GPU memory, a rolling windows technique is employed to manage input data that exceeds these thresholds. This method involves segmenting each time series into rolling windows characterized by predetermined lengths and step sizes; i.e., a processed time series $\mathbf{X}_{q}$ is segmented and turned into a set $\{\left(x^{i}_{q_{1\dots w}}\right)\}_{i=1}^{N}$, where $w$ is the window size and $N$ is the number of windows. For a cleaner notation, we refer to the set as $\{\left(x^{i}_{1\dots w}\right)\}_{i=1}^{N}$. We drop $q$ in the notation from this point on, as all the input is now quantized.
Tokenization. Different tokenization schemes vary in how they treat numerical values. Several open-source LLMs, such as Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2023), utilize the SentencePiece Byte-Pair Encoding tokenizer Touvron et al. (2023), which segments numbers into individual digits. However, the GPT tokenizer tends to segment numbers into chunks that may not correspond directly with the individual digits Liu and Low (2023). For instance, the number 234595678 is segmented into chunks [234, 595, 678] and assigned token IDs [11727, 22754, 17458]. Empirical evidence suggests that this segmentation impedes the LLM’s ability to learn patterns in time series data Gruver et al. (2023). To make sure GPT tokenizes each digit separately, we adopt the approach introduced by Gruver et al. (2023), which inserts spaces between the digits in a number. Continuing with the running example:
@@
Figure 3 shows how different preprocessing steps affect the output of the model. Overall, we find that scaling reduces the number of tokenized digits and yields better results than not scaling. Moreover, GPT performs better with added space between digits, while Mistral does not. These results accord with the forecasting representation presented in Gruver et al. (2023).

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated an outstanding ability to learn natural language tasks implicitly, whether performing reading comprehension, text summarization, translation, or related tasks. Radford et al. (2019); Brown et al. (2020); Sanh et al. (2022); Chowdhery et al. (2023); Wei et al. (2022). Moreover, LLMs have shown tremendous promise for formal language generation, including code generation and synthesis Xu et al. (2022a); Austin et al. (2021); Chen et al. (2021), and in production beyond textual output, such as generating images and videos from natural language descriptions Saharia et al. (2022); Koh et al. (2023). Testing these models on new tasks and data modalities allows us to push the boundaries of LLMs and discover their value. In this paper, we present a thorough study of LLMs used for the challenging task of anomaly detection from time series data, asking the question Can LLMs become anomaly detectors for time series data? Here, LLMs are exposed to a new data type – time series – and are tasked with a detection task – different from the classification tasks in which they are known to excel at Howard and Ruder (2018).
@@
Figure 1: (left) F1 Score performances of different model types, compared to a moving average baseline. Each category represents a collection of models that fall under that group. For classic models, we consider ARIMA and Matrix Profiling; for Deep Learning (DL), we utilize AER and LSTM DT; for transformer anomaly detection models, we look at Anomaly Transformer; lastly, for the commercial category, we compare to MS Azure. (right) Illustration of Mistral forecasts on E-2 signal from the SMAP dataset. The deviation between the signals can help identify anomalous regions.
@@
Recently, Gruver et al. (2023) posited that large language models have an inherent auto-regressive feature which allows them to be effective forecasters. In the study, the authors fed a string representation of a time series sequence to a pretrained LLM. The LLM then generated the next expected values, treating time series forecasting as a next-token prediction task. A follow-up question ari QWERT]śes: Does LLMs’ auto-regressive nature allow them to take on more complex tasks, such as anomaly detection? State-of-the-art time series anomaly detection models using deep learning typically include a forecasting model as one of the steps in their process Hundman et al. (2018).
@@
Time series anomaly detection is a regular part of day-to-day industry operations. Identifying unusual patterns can be a cumbersome and difficult task, especially when massive amounts of signal must be analyzed. If LLMs are genuine anomaly detectors, and can be employed directly in zero-shot (without any additional training), they could serve as off-the-shelf anomaly detectors for users, lifting a considerable amount of this burden. Considering that training deep learning models is time-consuming, skipping this phase could make anomaly detection more efficent overall.
@@
In this paper, we present SigLLM, a framework for using LLMs to detect anomalies in time series data, with current interaction support for models hosted by OpenAI  [^1] https://platform.openai.com/docs/models and HuggingFace  [^2] https://huggingface.co/models. Our framework includes a signal-to-text representation component to convert time series data into LLM-ready input. Moreover, we present two distinct approaches to investigating our main question. First, Prompter is a simple and direct prompting method, which elicits LLMs to identify the parts of a sequence it thinks are anomalous. Second, Detector leverages LLMs’ ability to forecast time series to find anomalies, by using the residual between the original signal and the forecasted one.
@@
Our findings, captured in Figure 1(left), show that LLMs improve on a simple moving average baseline. Moreover, they outperform transformer-based models such as Anomaly Transformer Xu et al. (2022b). However, there is still a gap between classic and deep learning approaches and LLMs. Furthermore, between our two approaches, Detector is superior to Prompter, with an improvement of 135% in F1 Score, as the latter suffers from false positives. We highlight the potential of the Detector approach in Figure 1(right), which showcases an example of an LLM forecast. We can clearly see that the LLM forecast is substantially different from the original signal; this difference is attributed to the presence of anomalies.
@@
  • Propose a new application for LLMs–anomaly detection–and study their efficacy and efficiency for this task. We formalize a new task to present to LLMs; namely, time series anomaly detection in zero-shot.
  • Present the SigLLM framework with a time-series-to-text representation module and two novel methodologies for solving this task. We present SigLLM with a module to convert time series data into language-model-ready input through a series of reversible transformations. Moreover, we propose two distinct approaches for solving the problem: the Prompter pipeline and the Detector pipeline. As of this writing, our framework integrates propriety models such as GPT by OpenAI and open models provided on the HuggingFace transformers package.
  • Provide a comprehensive and thorough evaluation of LLM performance on this task. We conduct our experiments on two prominent LLM models – GPT-3.5-turbo and Mistral-7B-Instruct-v0.2 – and 11 datasets. We show that LLMs are able to find anomalies with an average F1 score of 0.525. Moreover, we compare SigLLM methods to 10 other existing methods including state-of-the-art models such as AER Wong et al. (2022).
  • Publish an open source software. Our code and datasets are publicly available on github: https://github.com/sintel-dev/sigllm.
@@ 2 Background and Related Work
Anomaly Detection Pipelines. Unsupervised machine learning-based anomaly detection pipelines generally follow the same sequence of steps, which roughly consist of pre-processing, modeling, and post-processing abstractions as presented in Figure 2 Alnegheimish et al. (2022). Pre-processing operations include scaling the time series into a specific range, while post-processing includes computing discrepancies between two sequences. A wide variety of models can be trained to learn the features of input data, including Long-Short-Term-Memory models (LSTMs) Hundman et al. (2018), AutoEncoders (AE) Malhotra et al. (2016), Variational AutoEncoders (VAE) Park et al. (2018), Generative Adversarial Networks (GANs) Geiger et al. (2020), and Transformers Xu et al. (2022b); Tuli et al. (2022). These models perform well on existing benchmarks such as Alnegheimish et al. (2024), surpassing the performance of statistical approaches such as ARIMA Box and Pierce (1970); Pena et al. (2013).
@@
Transformers for Time Series. Transformers can be used directly to reconstruct time series Tuli et al. (2022). Moreover, the attention mechanism can be leveraged to find anomalous sequences Xu et al. (2022b). Recently, more work has emerged adopting transformer-based models for forecasting purposes by pretraining large transformer models on a large corpus of time series data. ForecastPFN Dooley et al. (2023) pre-trains a basic encoder-decoder transformer with one multi-head attention layer and two feedforward layers on a synthetically generated time series dataset. Similarly, TimeGPT was pretrained on a large collection of publicly available time series datasets. Lag-Llama Rasul et al. (2023) is a decoder-only Llama-2 model pretrained on a large corpus of real time series data from diverse domains. Moreover, Chronos Ansari et al. (2024) adopts a T5 architecture, and parses time series data into text to pretrain their model. Most of these models were developed with the objective of creating a time series foundation model for time series forecasting.
@@
LLMs for Time Series. The past several months have seen considerable efforts toward LLM utilization for time series data. Given the parallels between predicting the next word in a sentence and predicting the next value in a time series, most of these efforts have focused on time series forecasting. One notable effort is LLMTime where Gruver et al. (2023) employ GPT Brown et al. (2020), and Llama-2 Touvron et al. (2023) models to forecast time series data. PromptCast Xue and Salim (2023) is a related work that translates a forecasting problem into a prompt, transforming forecasting into a question-answering task.
@@
Figure 3: Visualizing the output of large language models (GPT and Mistral) under different variations of the transformation process. Each row depicts the exchange-2_cpm_results signal from the AdEx dataset, where the x-axis shows the timestamp and the y-axis is the signal value. The first row indicates the ground truth anomalies present in the time series (highlighted in green). The remaining rows indicate whether scaling and inserting space between digits has occurred during the conversion from signal to text. The gray intervals highlight the anomalies detected under these conditions; thus, we would like to maximize the overlap between the green and gray intervals. Overall we find that “scaling + space” is the configuration that yields a better output for GPT; and “scaling + no space” is better for Mistral.
@@
Our Work. In this paper, we work strictly with LLMs that have been pre-trained on text, particularly a proprietary model using GPT-3.5 Brown et al. (2020) and an open source model using Mistral Jiang et al. (2023). Our main objective is to determine whether LLMs have the ability to directly uncover anomalies in time series data. Referring back to Figure 2, our methodology focuses on Step 2 onwards – primarily the inference phase. To our knowledge, there is no other work that utilizes large language models as zero-shot anomaly detectors for time series data. We explore two avenues for accomplishing this task: (a) Through the paradigm of prompt engineering; (b) By leveraging LLMs’ ability to forecast time series in zero-shot without any additional data or fine-tuning.
@@ 3 Time Series Representation
Time series data can take many different forms. In this paper, we define a univariate time series as $\mathbf{X}=(x_{1},x_{2},\dots,x_{T})$, where $x_{t}\in\mathbb{Z}_{\geq 0}$ is the value at time step $t$, and $T$ is the length of the series. To make a time series LLM-ready, we transform the univariate time series $\mathbf{X}$ into a sequence of values that is tokenized. We follow a sequence of reversible steps, beginning with scaling, quantization, and processing the time series into segments using rolling windows, and ending with tokenizing each window. We detail these steps below.
@@
Figure 4: Anomaly detection methods in the SigLLM framework. (a) Prompter: a prompt engineering approach to elicit large language models to identify parts of the input which are anomalies. (b) Detector: a forecasting approach to use large language models as forecasting methods. Detector then finds discrepencies between the original and forecasted signal, which indicate the presence of anomalies.
Other scaling methods, such as min-max scaling, can be utilized to achieve the same goal. However, reducing the set of possible values to a smaller range (e.g. [0, 1]), may cause a loss of information in the quantization step. On the other hand, increasing the range will mean there are more digits to tokenize. With our approach, we simply shift the range of the signal values, which allows us to reduce the number of individual digits that need to be tokenized while maintaining the original gaps between pairs of entries. Moreover, by projecting the values into a non-negative range, we eliminate the need for sign indicator “-/+” and save an additional token.
Quantization. Unlike the finite set of vocabulary words used to train LLMs (32k vocab tokens for Mistral [^3] The exact vocabulary size for GPT-3.5-turbo has not been released by OpenAI., the set of scaled time series values $x_{s_{t}}$ is infinite, and cannot be processed by language models. Therefore, time series that are to be used with LLMs are generally quantized Ansari et al. (2024); Gruver et al. (2023). We use the rounding method, as proposed in in Gruver et al. (2023). Because in some cases the number of decimal digits are redundant given a fixed precision, we round each value up to a predetermined number of decimals, and subsequently scale to an integer format to avoid wasting tokens on the decimal point. Hence, the input time series becomes $\mathbf{X}_{q}=(x_{q_{1}},x_{q_{2}},\dots,x_{q_{T}})$, where $x_{q_{t}}\in\mathbb{Z}_{\geq 0}$. Below is an example of this operation :
@@
Overall, we use 2 mapping functions: the scaling function noted $\mathcal{E}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ and the quantization function noted $\mathcal{Q}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{Z}_{\geq 0}$. Because both mapping functions are reversible up to a certain number of precision digits, we can always reconstruct the input time series: $\mathcal{E}^{-1}\left(\mathcal{Q}^{-1}\left(x_{q_{t}}\right)\right)\approx x_{t}$
Rolling windows. Because there is an upper limit on the context length input to LLMs (e.g., Mistral has an upper limit of 32k tokens and GPT-3.5-turbo has a limit of 16k tokens), and there are constraints on GPU memory, a rolling windows technique is employed to manage input data that exceeds these thresholds. This method involves segmenting each time series into rolling windows characterized by predetermined lengths and step sizes; i.e., a processed time series $\mathbf{X}_{q}$ is segmented and turned into a set $\{\left(x^{i}_{q_{1\dots w}}\right)\}_{i=1}^{N}$, where $w$ is the window size and $N$ is the number of windows. For a cleaner notation, we refer to the set as $\{\left(x^{i}_{1\dots w}\right)\}_{i=1}^{N}$. We drop $q$ in the notation from this point on, as all the input is now quantized.
Tokenization. Different tokenization schemes vary in how they treat numerical values. Several open-source LLMs, such as Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2023), utilize the SentencePiece Byte-Pair Encoding tokenizer Touvron et al. (2023), which segments numbers into individual digits. However, the GPT tokenizer tends to segment numbers into chunks that may not correspond directly with the individual digits Liu and Low (2023). For instance, the number 234595678 is segmented into chunks [234, 595, 678] and assigned token IDs [11727, 22754, 17458]. Empirical evidence suggests that this segmentation impedes the LLM’s ability to learn patterns in time series data Gruver et al. (2023). To make sure GPT tokenizes each digit separately, we adopt the approach introduced by Gruver et al. (2023), which inserts spaces between the digits in a number.
Continuing with the running example:
@@
Figure 3 shows how different preprocessing steps affect the output of the model. Overall, we find that scaling reduces the number of tokenized digits, and yields better results than not scaling. Moreover, GPT performs better with added space between digits, while Mistral does not. These results accord with the forecasting representation presented in Gruver et al. (2023).